OpenStudy (anonymous):

If g(x) = x2 - 5, find g(7)

OpenStudy (diyadiya):

Just replace x with 7

OpenStudy (hba):

just plugin 7

OpenStudy (diyadiya):

\[g(x)= x^2 - 5\] \[g(7)= (7)^2 -5\]